OpenStudy (hang254):

Are the two statements logically equivalent? Why or why not? (p ∧ ~q) and ~(p ∨ q)

OpenStudy (hang254):

@Hero

OpenStudy (hang254):

i say yes, because they are true

OpenStudy (phi):

logically equivalent means they give the same T or F when you put in the same p and q to do this problem you list all possible values for p and for q p q ~q p ^ ~q F F T F ^ T = F F T F F ^ F = F T F T T ^ T = T T T F T ^ F= F now do the same for ~(p v q)

OpenStudy (hang254):

T F T F

OpenStudy (phi):

The first line says: let p be F and q be F the ~q (not q) is the "flip" of q if q is F then ~q is T and vice versa then for p ^ ~q you put in F for p and T for ~q and do the "and" operation of course F and T is F you do that for each row.

OpenStudy (anonymous):

Also ~( p v q) can be translated to (~p v ~q)

OpenStudy (phi):

yes that is the table for ~(p v q) it does NOT match (p ∧ ~q) so the expressions are not logically equivalent as chaguanas points out (with a typo) ~(p v q) is logically equivalent to ~p ^ ~q and that is different from p ∧ ~q

OpenStudy (hang254):

so, you are saying that they are not logically equivalent

OpenStudy (phi):

you get different answers, so they are not equivalent if you tested ~(p v q) and ~p ^ ~q you would see these two are logically equivalent

OpenStudy (hang254):

alright, i can see what you are saying now.

OpenStudy (hang254):

Thanks phi !

OpenStudy (phi):

for ~(p v q) I get T F F F if you then compute (p ∧ ~q) = ~(p ∨ q) F = T F F = F T T = F F F = F T if the = operation gave you T for every choice the two expression would be equivalent